# Advancing Glitch Classification in Gravity Spy: Multi-view Fusion with Attention-based Machine Learning for Advanced LIGO’s Fourth Observing Run

### v1

@@ 2.2 Multi-view Fusion Strategy
As Gravity Spy generates a dataset of glitches with four different durations, it is crucial to have a method for incorporating glitch information of varying characteristic timescales. Multi-view fusion, a powerful strategy in machine learning, combines information from multiple sources to enhance the representation and performance of a model. Each view contributes unique and complementary insights, leading to improved accuracy, robustness, and generalization of the model huang2020fusion . As mentioned above, our Gravity Spy project serves as an example where a single image with four time windows enables models to extract features across diverse timescales. Typically, multi-view fusion can employ one of three fusion architectures with different strategies to combine data from various view images. We review those below: the performance of these three fusion strategies will be compared in the result section.
Figure 3: Illustration of three fusion strategies. (a) Early fusion; (b) Late fusion, and (c) Intermediate fusion
@@ 2.2.1 Early Fusion
[…]
Although it offers several advantages, it has some limitations that may impact its practicality, particularly for large-sized images. In previous classifiers glanzer2023data ; soni2021discovering ; zevin2017gravity ; bahaadini2017deep , the resizing of each view to $140\times 170$ was required to fit within the model’s limitations, leading to a loss of valuable information. Additionally, early fusion assumes equal contributions from all modalities to the task, which may not hold true in this specific project. For instance, Repeating Blips might only appear in the 4 s time window view and not in others, necessitating the model to assign greater weight to the 4 s window for accurate classification.
@@ 2.2.2 Late Fusion
Late fusion is a decision-level data fusion technique, where predictions or outputs from individual models trained on different views are combined at the final stage, as shown in Fig. 3b. This fusion process often involves aggregation methods, such as majority voting or averaging, to make the final decision. The advantage of late fusion is its flexibility in utilizing different models for each view, enabling the selection of the most suitable model for each input. Moreover, it handles missing views by excluding the corresponding model during the fusion process, ensuring model evaluation even when certain data sources are unavailable (though that was not in the case in this application). However, one drawback of late fusion is its potential to overlook cross-modal relationships, as fusion occurs at a higher decision level rather than at the earlier processing stages. Consequently, in our case, this might lead to suboptimal exploitation of valuable complementarity between different time-window views.
@@ 2.2.3 Intermediate Fusion
Intermediate fusion, also known as feature-level fusion, allows features from each view to be processed separately through different branches of the network, and then combines the resulting feature maps at an intermediate layer, as shown in Fig. 3c. This technique effectively addresses the limitations of both early fusion and late fusion. Specifically, by leveraging the complementarity of different views and capturing cross-modal relationships at a stage where features are still semantically meaningful, intermediate fusion enables the model to learn complex representations and potentially achieve better performance. One general issue with intermediate fusion is the handling of missing views. As the model relies on inputs from all views, the absence of one or more views can lead to the loss of crucial information, potentially affecting the model’s decision-making process. This problem is not a concern for this project, as each glitch always contains complete data from all four time windows during pre-processing.
@@ 2.3 Methods
The proposed classifier distinguishes itself from glitch classification through several key optimizations, addressing the limitations of previous models glanzer2023data ; soni2021discovering ; zevin2017gravity ; bahaadini2017deep . First, it adopts intermediate fusion, effectively combining four views to enhance its representation power by leveraging their complementary information. Second, to handle a larger input dimension, a deeper network is adopted, featuring the inception residual block szegedy2017inception to mitigate the vanishing gradient problem and to extract more discriminative features. Additionally, an attention module is introduced at a later stage, providing valuable insights into the model’s focus on specific views during decision-making. Last, the incorporation of smooth labeling during training effectively addresses the noisy label issue. The architecture of the proposed classifier is shown in Fig. 4, and detailed explanations of these optimizations are presented next.
Figure 4: Architecture of the proposed machine learning classifier used in O4. Further elaboration on the hyperparameters can be found in Section 2.3.
@@ 2.3.1 Inception Residual Block
The inception residual block combines the ideas of the inception module and the residual block proposed in Szegedy et al. szegedy2017inception . As shown in Fig. 4, it combines these concepts by applying the inception module within a residual block structure. Specifically, the block consists of four parallel convolutional branches with filters of different sizes ($1\times 1$, $1\times 1$, $3\times 3$, $5\times 5$). The outputs of these branches are then concatenated, and the resulting feature map is combined with the original input after a $1\times 1$ convolutional layer using a skip connection, followed by a rectified linear unit (ReLU) activation function. Particularly, the use of parallel convolutional filters of different sizes allows the block to capture features at multiple scales, improving the model’s ability to recognize patterns of varying complexities szegedy2015going . Moreover, the incorporation of skip connections facilitates the training of deep networks, effectively mitigating the vanishing gradient problem he2016deep .
@@ 2.3.2 Attention Module
To assess the varying importance of each view in the final glitch classification, we introduce an attention module ilse2018attention $\mathcal{M}_{\mathrm{Att}}$ to the latent vectors of four inputs, denoted as $z_{i}\in\mathbf{R}^{1\times M},i=\{1,2,3,4\}$, extracted after the fully-connected layer. The attention weight $\alpha_{i}\in\mathbf{R},i=\{1,2,3,4\}$ calculated for each latent vector $z_{i}$ quantifies the significance of each view in making the glitch classification, with all $\alpha_{i}$’s summing up to 1. The attention module is defined as:
@@
$\alpha_{i}=\frac{\exp\{w^{T}[\tanh(Vz_{i}^{T})\odot\sigma(Uz_{i}^{T})]\}}{\sum_{j=1}^{4}\exp\{w^{T}[\tanh(Vz_{j}^{T})\odot\sigma(Uz_{j}^{T})]\}}.$ (2)
The trainable parameters $V\in\mathbf{R}^{L\times M}$ and $w\in\mathbf{R}^{L\times 1}$ are associated with the attention mechanism. Here, $M$ represents the dimension of latent vectors (set to $M=128$ in this paper) while $L$ signifies the one dimension of the matrices for transforming the latent vectors (set to $L=64$). The $V$ matrix is instrumental in identifying similarities among different views. The non-linear hyperbolic tangent function $\tanh(\cdot)$ includes both positive and negative values in the gradient flow, but the concern is that it is almost linear in the $[-1,1]$ range. To handle complex data patterns effectively, we leverage the gated attention mechanism ilse2018attention , which combines $\tanh(\cdot)$ with the element-wise sigmoid function $\sigma(\cdot)$. The sigmoid function helps adjust the relevance of information, while the element-wise multiplication $\odot$ facilitates the attention process. The gated attention mechanism introduces additional parameters $U\in\mathbf{R}^{L\times M}$ to serve as control gates, mitigating the linearity issue present in $\tanh(\cdot)$. The attention module, widely used in multiple-instance learning tasks ilse2018attention ; wu2021combining ; wu2023smooth ; lopez2022deep , excels in identifying key instances within a bag, i.e., a collection of instances.
In this paper, we assess the impact of the attention module on the model’s performance through comparative experiments conducted with and without the incorporation of attention modules.
@@ 2.3.3 Label Smoothing
As mentioned in Sec. 2.1, the ground truth labels for the Gravity Spy dataset are generated by the previous classifier and volunteers. Despite filtering the dataset to only use high-confidence predictions, mislabeled glitches can still be present, potentially confusing the model. To address this problem, label smoothing muller2019does , a valuable regularization technique in machine learning, is applied during model training to enhance generalization and prevent overconfidence in predictions, especially when dealing with noisy or uncertain data. This technique introduces a small amount of uncertainty into the training labels. Instead of assigning a hard 1 to the correct class and 0 to other classes, label smoothing assigns a smoothed value, denoted by $q$, slightly less than 1 to the correct class while distributing the remaining confidence uniformly across the other classes. Mathematically, it can be defined as:
@@
In addition, to address the issue of class imbalance (as shown in Table 1), we apply different weights $\gamma_{i}$ to each class $i=\{1,2,..C\}$ in the loss function during training. These weights are inversely proportional to the number of instances in each glitch class and sum up to 1, ensuring a balanced learning process. Therefore, the loss function of the proposed model is defined as
$\mathcal{L}(x,y)=-\frac{1}{C}\sum_{i=1}^{C}\gamma_{i}q_{i}\log(\hat{y_{i}}),$ (4)
@@ 2.3.4 Architecture of the Classifier
The overall architecture of the proposed classifier is shown in Fig. 4. Each of the four branches shares a common architecture. Within each branch, a pre-convolutional block is followed by three instances of a combination of a convolutional block and an inception residual block (Incep-Res Block) szegedy2017inception , along with a post-convolutional block. These stages effectively extract features from each glitch view. Specifically, all convolutional blocks (i.e., Pre-Conv, Conv and post-Conv blocks) consist of a convolutional layer, a ReLu activation function, and batch normalization (BN) layer and a max pooling (Max Pool) layer. Subsequently, a flatten layer and a fully-connected (FC) layer process and reduce the dimensionality of the feature vectors. The attention module ilse2018attention then takes these vectors as input, generating a single weighted averaged feature vector, which is then fed into the final FC layer. The softmax activation function yields the probabilities of the 23 glitch classes. To enhance the model’s generalization and prevent overfitting, we incorporate an $L2$ regularizer cortes2012l2 for each convolutional layer and introduce dropout srivastava2014dropout after each FC layer. For optimization, we use the Adam optimizer kingma2014adam , with the learning rate starting at $2\times 10^{-5}$. This choice of optimizer helps to efficiently update the model’s parameters during training. To address the potential issue of confidence overfitting, where the model may become overconfident in its predictions on the training data, we employ early stopping with a patience of 10, i.e., if the model doesn’t improve its accuracy on the validation set within 10 rounds of training, we stop training to prevent potential overconfidence.
@@ 2.4 Statistical Analysis
To ensure the robustness of the model, five independent runs for all experiments are conducted. For each experiment, we calculate the overall accuracy, precision, F1 score, and the area under the receiver operating characteristic curve (AUC) on the held-out test set, where the mean and standard deviation values for each metric are reported. To assess the significance of the differences between different model results, we perform a paired t-test kim2015t , considering the comparison statistically significant if the $p$-value is less than 0.05 ($p<0.05$). Furthermore, to calibrate the performance of the model, a reliability plot brocker2007increasing is employed, offering a visual representation of the model’s calibration across varying confidence levels. Specifically, this plot juxtaposes predicted probabilities against empirical frequencies of positive outcomes, providing insights into the model’s accuracy in conveying uncertainty. Additionally, calibration accuracy is quantified through the expected calibration error (ECE) guo2017calibration , a statistical metric that measures the average deviation between predicted probabilities and the true likelihood of event occurrence.
@@ 3.1 Experimental Design
In this study, we conduct various ablation studies to assess the model’s performance. First, the effectiveness of different fusion strategies is compared, as shown in Fig. 3. Specifically, the early fusion involves merging four view images and resizing the merged input to a size of $448\times 448\times 3$ before feeding it into the model. The late fusion uses each branch to individually produce probabilities for each class, and the final classification is determined by averaging the probabilities obtained from the four view images. The proposed classifier employs intermediate fusion. Second, we evaluate different values of the smoothing parameter $\beta$ between 0 and 1 with an increment of 0.1, allowing us to identify the optimal value for $\beta$. Third, we compare the proposed classifier with the attention module and a variant without the attention module (i.e., max pooling and mean pooling) to demonstrate the effectiveness of the attention mechanism. Fourth, we evaluate the performance of our proposed classifier by comparing it with the previous Gravity Spy model architecture soni2021discovering ; glanzer2023data . Both models are trained from scratch using the same training dataset, and their performance is assessed on the same test set. Finally, to better understand the learned feature representations, we employ t-distributed stochastic neighbor embedding (t-SNE) plots to reduce the dimensionality of feature vectors obtained after the attention module from 128 to 2, offering insightful visualizations for model comparisons in a two-dimensional space. The experiments and analyses are conducted using TensorFlow 2.11 on two GPUs (NVIDIA Quadro RTX 8000) and scikit-learn, respectively, in Python 3.8. The codes are available via the GitHub of Gravity Spy.
@@ 3.2 Comparison of Three Fusion Strategies
Table 2 demonstrates the model performance of three fusion strategies on the same test set. The proposed intermediate fusion model achieves the best result with an overall accuracy of 0.941, precision of 0.950, F1 of 0.931 and AUC of 0.965, followed by the late fusion and the early fusion, although the difference between the latter two is not statistically significant. Moreover, the results exhibit robustness, as indicated by the minimal standard deviations derived from 5 independent experiment runs.
Table 2: Model evaluations (the overall mean ± standard deviation across 5 independent runs) for three fusion strategies (i.e., early fusion, late fusion and intermediate fusion) on the test set. Best results are highlighted in bold, and * indicates that the comparison is statistically significant ($p<0.05$).
Fusion
Strategies
Early Fusion
(Fig. 3a)
Late Fusion
(Fig. 3b)
Intermediate Fusion
(Proposed Fig. 4)
@@
3.3 Effects of Smoothing Levels
Figure 5 presents a comparative analysis of model performance across various smoothing levels of $\beta$ in Eq. (3). At the outset, for $\beta=0$, no label smoothing yields a model accuracy of 0.921, precision of 0.915, F1 of 0.908 and AUC of 0.946. With an increase in $\beta$, the model’s performance exhibits an initial ascent followed by a subsequent decline. When $\beta=0.3$, the model achieves the optimal performance across all metrics. Additionally, the results indicate that, with the implementation of label smoothing, for $\beta\leqslant 0.5$, the model consistently outperforms the one without label smoothing ($\beta=0$)..
Figure 5: The evaluation of model performance (Accuracy, Precision, F1, and AUC scores) on the test set presented across varying smoothing levels ($\beta$) ranging from 0 to 1 in increments of 0.1. Here, $\beta=0$ represents no smoothing.

### v2

@@ 2.2 Multi-view fusion strategy
As Gravity Spy generates a dataset of glitches with four different durations, it is crucial to have a method for incorporating glitch information of varying characteristic timescales. Multi-view fusion, a powerful strategy in machine learning, combines information from multiple sources to enhance the representation and performance of a model. Each view contributes unique and complementary insights, leading to improved accuracy, robustness, and generalization of the model [33]. As mentioned above, our Gravity Spy project serves as an example where a single image with four time windows enables models to extract features across diverse timescales. Typically, multi-view fusion can employ one of three fusion architectures with different strategies to combine data from various view images. We review those below: the performance of these three fusion strategies will be compared in Section 3.
Figure 3: Illustration of three fusion strategies: (a) early fusion; (b) late fusion, and (c) intermediate fusion. CNN: convolutional neural network; NNs: neural networks.
@@ 2.2.1 Early fusion
Early fusion takes place at the input level, where features from each view are combined in the input layer of the network before being fed into one single model. By merging data from multiple views at this stage, the model has access to a more comprehensive representation of the data. This approach was employed in previous Gravity Spy classifiers [24, 17, 32], which constructed a $2h\times 2w$ matrix by juxtaposing four $h\times w$ view images, as illustrated in Figure 3 (a).
Although it offers several advantages, early fusion has some limitations that may impact its practicality, particularly for large-sized images. In previous classifiers [24, 17, 32], the resizing of each view to $140\times 170$ was required to fit within the model’s limitations, leading to a loss of valuable information. Additionally, early fusion assumes equal contributions to the task from all modalities, which may not hold true in this specific project. For instance, Repeating Blips [25] might only appear in the $4~\mathrm{s}$ time-window view and not in others, necessitating the model to assign greater weight to the $4~\mathrm{s}$ window for accurate classification.
@@ 2.2.2 Late fusion
Late fusion is a decision-level data-fusion technique, where predictions or outputs from individual models trained on different views are combined at the final stage, as shown in Figure 3 (b). This fusion process often involves aggregation methods, such as majority voting or averaging, to make the final decision. The advantage of late fusion is its flexibility in using different models for each view, enabling the selection of the most suitable model for each input. However, one drawback of late fusion is its potential to overlook cross-modal relationships, as fusion occurs at a higher decision level rather than at the earlier processing stages [41, 42]. Consequently, in our case, this might lead to suboptimal exploitation of valuable complementarity between different time-window views.
@@ 2.2.3 Intermediate fusion
Intermediate fusion, also known as feature-level fusion, allows features from each view to be processed separately through different branches of the network, and then combines the resulting feature maps at an intermediate layer, as shown in Figure 3 (c). This technique effectively addresses the limitations of both early fusion and late fusion. Specifically, by leveraging the complementarity of different views and capturing cross-modal relationships at a stage where features are still semantically meaningful, intermediate fusion enables the model to learn complex representations and potentially achieve better performance.
@@ 2.3 Methods
The proposed classifier distinguishes itself from the past Gravity Spy glitch classifier through several key optimizations, addressing the limitations of previous models [25, 24, 17, 32]. First, it adopts intermediate fusion, effectively combining four views to enhance its representation power by leveraging their complementary information. Second, to handle a larger input dimension (higher resolution images), a deeper network is adopted, featuring the inception residual block [43] to mitigate the vanishing gradient problem and to extract more discriminative features. Additionally, an attention module is introduced at a later stage, providing valuable insights into the model’s focus on specific views during decision-making. Last, the incorporation of smooth labeling during training effectively addresses the noisy label issue. The architecture of the proposed classifier is shown in Figure 4, and detailed explanations of these optimizations are presented next.
Figure 4: Architecture of the proposed machine-learning classifier used in O4. Further elaboration on the hyperparameters can be found in Section 2.3. The $\oplus$ represents element-wise addition, while the $\otimes$ represents multiplication by attention weights in Eq. (1). FC: fully connected; BN: batch normalized.
@@ 2.3.1 Inception residual block
The inception residual block combines the ideas of the inception module and the residual block proposed in Szegedy et al. [43]. As shown in Figure 4, it combines these concepts by applying the inception module within a residual block structure. Specifically, the block consists of four parallel convolutional branches with filters of different sizes ($\times 1, 1×1$$\times 1$, $3\times 3$, $5\times 5$). There is a $1\times 1$ convolution, serving as a dimensionality-reduction layer, before the computationally expensive $3\times 3$ and $5\times 5$ convolutions that reduces the computational cost while preserving the network’s capability to learn complex patterns. The outputs of the four branches are concatenated, and the resulting feature map is combined with the original input after a $1\times 1$ convolutional layer using a skip connection, followed by a rectified linear unit (ReLU) activation function. The use of parallel convolutional filters of different sizes allows the block to capture features at multiple scales, improving the model’s ability to recognize patterns of varying complexities [44]. This is useful in our case, as different glitch classes have time–frequency features at a variety of scales. Moreover, the incorporation of skip connections facilitates the training of deep networks, effectively mitigating the vanishing gradient problem [45].
@@ 2.3.2 Attention module
To assess the varying importance of each view in the final glitch classification, we introduce an attention module $\mathcal{M}_{\mathrm{Att}}$ [46] to the latent vectors of four inputs, denoted as $z_{i}\in\mathbb{R}^{1\times M},i=\{1,2,3,4\}$, extracted after the fully-connected layer. We set the dimension of the latent vectors to $M=128$ for this model. The attention weight $\alpha_{i}\in\mathbb{R},i=\{1,2,3,4\}$ calculated for each latent vector $z_{i}$ quantifies the significance of each view in making the glitch classification, with all weights summing up to $1$. The attention module is defined as
@@
$\alpha_{i}=\frac{\exp\{w^{\mathrm{T}}[\tanh(Vz_{i}^{\mathrm{T}})\odot\sigma(Uz_{i}^{\mathrm{T}})]\}}{\sum_{j=1}^{4}\exp\{w^{\mathrm{T}}[\tanh(Vz_{j}^{\mathrm{T}})\odot\sigma(Uz_{j}^{\mathrm{T}})]\}}.$ (2)
The trainable parameters $V\in\mathbb{R}^{L\times M}$ and $w\in\mathbb{R}^{L\times 1}$ are associated with the attention mechanism. Here, $L$ signifies the one dimension of the matrices for transforming the latent vectors (set to $L=64$). The $V$ matrix is instrumental in identifying similarities among different views. The non-linear hyperbolic tangent function $\tanh(\cdot)$ includes both positive and negative values in the gradient flow, but it is almost linear in the [1,1][-1,1] range. To handle complex data patterns effectively, we leverage the gated attention mechanism [46], which combines $\tanh(\cdot)$ with the element-wise sigmoid function $\sigma(\cdot)$. The sigmoid function helps adjust the relevance of information, while the element-wise multiplication $\odot$ facilitates the attention process. The gated attention mechanism introduces additional parameters $U\in\mathbb{R}^{L\times M}$ to serve as control gates, mitigating the linearity issue present in $\tanh(\cdot)$. The attention module, widely used in multiple-instance learning tasks [46, 47, 48, 49], excels in identifying key instances within a bag, i.e., a collection of instances.
In this paper, we evaluate the impact of the attention module on the model’s performance by conducting comparative experiments both with and without the incorporation of attention. For the experiments without the attention module, we employed two approaches: averaging the latent vector (i.e., mean pooling) and maximizing the latent vector (i.e., max pooling).
@@ 2.3.3 Label smoothing
As mentioned in Section 2.1, the ground-truth labels for the Gravity Spy dataset are generated by the previous classifier and volunteers. Despite filtering the dataset to only use high-confidence predictions, mislabeled glitches can still be present, potentially confusing the model. To address this problem, label smoothing [50], a valuable regularization technique in machine learning, is applied during model training to enhance generalization and prevent overconfidence in predictions, especially when dealing with noisy or uncertain data. This technique introduces a small amount of uncertainty into the training labels. Instead of assigning a hard $1$ to the correct class and 0 to other classes, label smoothing assigns a smoothed value, denoted by $q$, slightly less than $1$ to the correct class while distributing the remaining confidence uniformly across the other classes. Mathematically, it can be defined as:
@@
In addition, to address the issue of class imbalance (as shown in Table 1), we apply different weights $\gamma_{i}$ to each class $i=\{1,2,\ldots,C\}$ in the loss function during training. These weights are inversely proportional to the number of instances in each glitch class and sum up to $1$, ensuring a balanced learning process. Therefore, the loss function of the proposed model is defined as
[…]
@@ 2.3.4 Architecture of the classifier
The overall architecture of the proposed classifier is shown in Figure 4. Each of the four branches shares a common architecture. Within each branch, a pre-convolutional block is followed by three instances of a combination of a convolutional block and an inception residual block (Incep-Res Block) [43], along with a post-convolutional block. These stages effectively extract features from each glitch view. Specifically, all convolutional blocks (i.e., Pre-Conv, Conv and Post-Conv blocks) consist of a convolutional layer, a ReLu activation function, and batch normalization (BN) layer and a max pooling (Max Pool) layer. Subsequently, a flatten layer and a fully-connected (FC) layer reduce the dimensionality of the feature vectors. The attention module [46] then takes these vectors as input, generating a single weighted-averaged feature vector, which is then fed into the final fully-connected layer. The softmax activation function yields the probabilities of the $23$ glitch classes. To enhance the model’s generalization and prevent overfitting, we incorporate an $L2$ regularizer [51] for each convolutional layer and introduce dropout [52] after each fully-connected layer.
For optimization, we use the Adam optimizer [53], with the learning rate starting at $2\times 10^{-5}$. This choice of optimizer helps to efficiently update the model’s parameters during training. To address the potential issue of confidence overfitting, where the model may become overconfident in its predictions on the training data, we employ early stopping with a patience of $10$, i.e., if the model does not improve its accuracy on the validation set within $10$ training epochs, we stop training to prevent potential overconfidence.
@@ 2.4 Statistical analysis
To ensure the robustness of the model, five independent runs for all experiments are conducted. Each of the five runs has different random initializations of the network weights and uses different random mini-batch orderings during training, while keeping the same architecture and hyperparameters. For each experiment, we calculate the overall accuracy, precision, F1 score (harmonic mean of the precision and recall), and the area under the receiver operating characteristic curve (AUC) on the held-out test set, where the mean and standard deviation values for each metric are reported. To assess the significance of the differences between different model results, we perform a paired $t$-test [54] by calculating the differences in scores for each pair, followed by computing the mean and standard deviation of these differences, and determining significance based on a fiducial p-value ($p<0.05$). Furthermore, we employ a reliability plot [55] to evaluate model calibration across different confidence levels. This plot compares predicted probabilities with observed outcomes by binning the probabilities and calculating the average predicted values and observed frequencies for each bin. By juxtaposing these elements, the plot reveals the model’s accuracy in representing uncertainty. In other words, for a well-calibrated model, if it predicts a set of glitches with a $40\%$ probability of being a Blip, the actual frequency of Blips in that set should also be approximately $40\%$. Additionally, calibration accuracy is quantified through the expected calibration error (ECE) [34], a statistical metric that measures the average deviation between predicted probabilities and the true likelihood of event occurrence.
@@ 3.1 Experimental design
In this study, we conduct various ablation studies to assess the model’s performance. First, the effectiveness of different fusion strategies is compared, as shown in Figure 3. Second, we evaluate different values of the smoothing parameter $\beta$ between 0 and $1$ with an increment of $0.1$, allowing us to identify an optimal value for $\beta$. Third, we compare the proposed classifier with the attention module and a variant without the attention module (i.e., max pooling and mean pooling) to demonstrate the effectiveness of the attention mechanism. Fourth, we evaluate the performance of our proposed classifier by comparing it with the previous Gravity Spy model architecture [23, 24, 25]. Both models are trained from scratch using the same training dataset, and their performance is assessed on the same test set. Training and testing on the same dataset ensures a fair comparison; these have varied in previous studies (along with the list of glitch classes), so it would not be fair to directly compare statistics to those from other works. Finally, to better understand the learned feature representations, we employ $t$-distributed stochastic neighbor embedding ($t$-SNE) plots [56] to reduce the dimensionality of feature vectors obtained after the attention module from $128$ to $2$, offering visualizations for model comparisons in a two-dimensional space. Here, for each experiment, we maintain all other experimental conditions at their optimal settings to ensure a fair comparison.
The experiments and analyses are conducted using TensorFlow 2.11 on two GPUs (NVIDIA Quadro RTX 8000) and scikit-learn, respectively, in Python 3.8. The codes are available via the GitHub of Gravity Spy.
@@ 3.2 Comparison of three fusion strategies
Table 2 demonstrates the model performance of three fusion strategies on the same test set. The proposed intermediate fusion model achieves the best result with an overall accuracy of $0.941$, precision of $0.950$, F1 of $0.931$ and AUC of $0.965$, followed by the late fusion and the early fusion, although the difference between the latter two is not statistically significant. Moreover, the results exhibit robustness, as indicated by the minimal standard deviations derived from five independent experiment runs.
Table 2: Model evaluations (the overall mean $\pm$ standard deviation across five independent runs) for three fusion strategies (early fusion, late fusion and intermediate fusion) on the O3 test set. Best results are highlighted in bold, and an asterisk (*) indicates that the comparison is statistically significant given our chosen threshold ($p<0.05$).
Fusion strategy
Performance metric Early fusion Late fusion Intermediate fusion
(Figure 3 (a)) (Figure 3 (b)) (proposed; Figure 4)
@@
3.3 Effects of smoothing labels
Figure 5 presents a comparative analysis of model performance across various smoothing levels of $\beta$ in Eq. (3). At the outset, for $\beta=0$, no label smoothing yields a model accuracy of $0.921$, precision of $0.915$, F1 of $0.908$ and AUC of $0.946$. With an increase in $\beta$, the model’s performance exhibits an initial ascent followed by a subsequent decline. When $\beta=0.3$, the model achieves the optimal performance across all metrics. Additionally, the results indicate that, with the implementation of label smoothing, for $\beta\leqslant 0.5$, the model consistently outperforms the one without label smoothing ($\beta=0$).
Figure 5: The evaluation of model performance (accuracy, precision, F1, and AUC scores) on the O3 test set presented across varying smoothing levels ($\beta$) ranging from 0 to 11 in increments of $0.1$. The results represent the average of five experimental runs, with standard deviations indicated by error bars. Here, $\beta=0$ represents no smoothing.